OpenStudy (anonymous):

When under equal tension, the frequency of a vibrating string from a piano varies inversely with the string length. If a string that is 375 millimeters in length vibrates at a frequency of 462 cycles a second, at what frequency will a 620 millimetre string vibrate?

OpenStudy (anonymous):

Pleaseee help mee

OpenStudy (kash_thesmartguy):

\[\frac{ 462 }{ 375 }=\frac{ ? }{ 620 }\]"?" resembles your answer. To get to that, you will have to divide 462 by 375 which is 1.23 Then you have to multiply 620 by 1.23 So, \[6.20*1.23=YourAnswer\]

OpenStudy (kash_thesmartguy):

Welcome to OpenStudy! Here, you can enjoy by learning. Here is the legendary tutorial for our legendary OpenStudiers: http://openstudy.com/study#/updates/5525c0fce4b030ceb3f357e8 Created by: @Kash_TheSmartGuy

OpenStudy (anonymous):

The answer would be 7.626?

OpenStudy (anonymous):

@Kash_TheSmartGuy

OpenStudy (anonymous):

@Kash_TheSmartGuy

OpenStudy (kash_thesmartguy):

Yes, the answer will be 7.626, good job!

OpenStudy (kash_thesmartguy):

@Tht1dudekk